Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

OG #217 question

Expert replies
by jamesk486 » Thu Apr 19, 2007 6:42 am
A certain junior class has 1,000 students and a certain senior class has 800 students. Among these students, there are 60 sibling pairs, each consisting of 1 junior and 1 senior. If a student is to be selected at random from each class, what is the probabilty that the 2 students selected will be a sibling pair?
A. 30/40,000
B. 1/3600
C. 9/2000
D. 1/60
E. 1/15

why is it (60/1000)(1/800)?
60/1000= a junior who is a member of a sibling pair, but shouldn't it be 30? Bc 30 should be juniors and 30 should be seniors?
And then why not (60/800) or (30/800)?
i solved it like (30/1000)*(30/800)
Join the discussion
Source: — Problem Solving |

by Cybermusings » Thu Apr 19, 2007 8:13 am
Ans Choice A here is 3/40,000 and not 30/40,000

Probability that out of 1000 junior students, we pick from a student from a sibling pair = 30/1000

Probability that out of the 800 senior students we pick a student who is the other half of the same sibling pair =1/800

So 60/1000 * 1/800 = 6/100*1/800 = 3/40,000 = 3/40,000

You have to keep in mind that the second selection is made after the first selection is already through and hence we select 1/800 and not 60/800
Join the discussion

by Scott@TargetTestPrep » Mon Nov 12, 2018 7:26 am
jamesk486 wrote:A certain junior class has 1,000 students and a certain senior class has 800 students. Among these students, there are 60 sibling pairs, each consisting of 1 junior and 1 senior. If a student is to be selected at random from each class, what is the probabilty that the 2 students selected will be a sibling pair?
A. 30/40,000
B. 1/3600
C. 9/2000
D. 1/60
E. 1/15
The probability of selecting any one sibling from the 60 sibling pairs in the junior class is 60/1000. Once that person is selected, the probability of selecting his or her sibling from the senior class is 1/800; thus, the probability of a selecting a sibling pair is:

60/1000 x 1/800 = 3/50 x 1/800 = 3/40000

We would obtain the same answer if we made our first selection from the senior class: The probability of selecting any one sibling from the 60 sibling pairs in the senior class is 60/800. Once that person is selected, the probability of selecting his or her sibling from the junior class is 1/1000; thus, the probability of a selecting a sibling pair is also:

60/800 x 1/1000 = 3/40 x 1/1000 = 3/40000

Answer: 3/40000

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion

by Brent@GMATPrepNow » Mon Nov 12, 2018 8:36 am
jamesk486 wrote:A certain junior class has 1,000 students and a certain senior class has 800 students. Among these students, there are 60 sibling pairs, each consisting of 1 junior and 1 senior. If a student is to be selected at random from each class, what is the probabilty that the 2 students selected will be a sibling pair?
A. 30/40,000
B. 1/3600
C. 9/2000
D. 1/60
E. 1/15
P(selecting a sibling pair) = P(select a junior with a sibling AND select the senior who is that junior's sibling)
= P(select a junior with a sibling) x P(select the senior who is that junior's sibling[/u])
= 60/1000 x 1/800
= 60/800,000
= 3/40,000
= A

Note: P(select a junior with a sibling) = 60/1000, because 60 of the 1000 juniors have a sibling who is a senior.
P(select a senior who is that junior's sibling) = 1/800, because there are 800 senior's and only 1 of them is the sibling of the selected junior.

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by fskilnik@GMATH » Mon Nov 12, 2018 1:21 pm
jamesk486 wrote:A certain junior class has 1,000 students and a certain senior class has 800 students. Among these students, there are 60 sibling pairs, each consisting of 1 junior and 1 senior. If a student is to be selected at random from each class, what is the probabilty that the 2 students selected will be a sibling pair?
A. 3/40,000
B. 1/3600
C. 9/2000
D. 1/60
E. 1/15
$$\left( J \right)\,\,{\text{Junior}}\,\,:\,\,\,1000\,\,{\text{students}}\,\,,\,\,\,{J_1}\,,\,{J_2},\,\, \ldots \,\,,\,\,{J_{60}}\,\,{\text{among}}\,\,{\text{them}}\,\,\,\left( {{\text{the}}\,\,{\text{ones}}\,\,{\text{with}}\,\,{\text{senior}}\,\,{\text{siblings}}} \right)$$
$$\left( S \right)\,\,{\text{Senior}}\,\,:\,\,\,800\,\,{\text{students}}\,\,,\,\,\,{S_1}\,,\,{S_2},\,\, \ldots \,\,,\,\,{S_{60}}\,\,{\text{among}}\,\,{\text{them}}\,\,\,\left( {{\text{the}}\,\,{\text{ones}}\,\,{\text{with}}\,\,{\text{junior}}\,\,{\text{siblings}}} \right)$$
$$\left( {{J_1}\,,\,{S_1}} \right)\,\,;\,\,\left( {{J_2}\,,\,{S_2}} \right)\,\,;\,\, \ldots \,\,;\,\,\left( {{J_{60}}\,,\,{S_{60}}} \right)\,\,\,:\,\,\,{\text{pairs}}\,\,{\text{of}}\,\,{\text{siblings}}\,$$
$$? = P\left( {{\text{pair}}\,{\text{of}}\,{\text{siblings}}\,,\,\,{\text{in}}\,{\text{one}}\,J\,{\text{and}}\,{\text{one}}\,S\,{\text{extraction}}} \right)$$
$${\text{Total}}\,\,:\,\,\,1000 \cdot 800\,\,\,{\text{equiprobable}}\,\,\,{\text{selections}}\,\,\,\left[ {\left( {{J_m},{S_n}} \right)\,\,,\,\,\,{\text{where}}\,\,1 \leqslant m \leqslant 1000\,\,{\text{and}}\,\,\,1 \leqslant n \leqslant 800} \right]$$
$${\text{Favorable}}\,\,:\,\,60\,\,\,\,\,\left[ {\left( {{J_k},{S_k}} \right)\,\,,\,\,\,{\text{where}}\,\,1 \leqslant k \leqslant 60} \right]\,\,\,\,$$
$$? = \frac{{60}}{{1000 \cdot 800}} = \underleftrightarrow {\frac{{4 \cdot 15}}{{1000 \cdot 4 \cdot 200}}} = \frac{{4 \cdot 3 \cdot 5}}{{1000 \cdot 4 \cdot 5 \cdot 40}} = \frac{3}{{40 \cdot 1000}}$$


This solution follows the notations and rationale taught in the GMATH method.

Regards,
Fabio.
Fabio Skilnik :: GMATH method creator ( Math for the GMAT)
English-speakers :: https://www.gmath.net
Portuguese-speakers :: https://www.gmath.com.br
Join the discussion